OpenStudy (anonymous):

A large planetarium in the shape of a dome is being built. When it is complete, the base of the dome will have a circumference of 87 kilometers. What area of land was required for this planetarium?

OpenStudy (anonymous):

circumference = 2pi r 87 km = 2pi r r= 87/ (2pi) Now that you have r, find area using the equation Area = pi r^2

OpenStudy (anonymous):

so I divide 87 by 2??

OpenStudy (anonymous):

\[\frac{ 87 }{ 2\pi }\]

OpenStudy (anonymous):

Is it 136??

OpenStudy (anonymous):

136.66

OpenStudy (anonymous):

That's wrong

OpenStudy (anonymous):

No that's not the answer. You need to find area. That's only radius again, area = pi r^2